Physics Free Fall Motion

Option One

Materials: Calculator

  1. In the table below, you are given the height from which various objects fell and the time it took for those objects to fall on Earth, the moon, and Mars. Use the data to calculate the acceleration due to gravitational force at each location.

     EarthMoonMars
    Distance (m)2.253.084.044.574.683.55
    Time (s)0.6800.7932.202.261.581.38
    Acceleration due to gravity (m/)





    Average g value for the location (m/)


  2. What to submit:

    1. Your completed table with calculations for acceleration due to gravitational force and the average value.

    2. All work shown for your calculations.
Answer #1

SOLUTION :


s = 0 + 1/2 g t^2

=> g = 2s / t^2 


=> g = 


Earth :


g earth1 = 2*2.25 / (0.680)^2 = 9.732 m/s^2


g earth2 = 2*3.08/0.793^2 = 9.796 m/s^2


avg. g earth = (9.732 + 9.796) / 2 = 9.764 m/s^2


Moon :


g moon1 = 2*4.04 / 2.20^2 = 1.669 m/s^2


G moon2 = 2*4.57 / 2.26^2 = 1/789 m/s^2


avg. g moon = (1.669 + 1.789) / 2 = 1.729 m/s^2 


Mars :


g mars1 = 2*4.68 / 1.58^2 = 3.749 m/s^2


g mars2 = 2*3.55 / 1.38^2 = 3.728 m/s^2


avg. g mars = (3.749 + 3.728 ) / 2 = 3.739 m/s^2 



Accordingly, table will be as under :



Particulars / Planet                                        Earth                 Moon                    Mars


Distance  (m)                                           2.25     3.08        4.04     4.57         4.68     3.55


Time (s)                                                   0.68     0.793       2.20    2.26         1.58     1.38


Acceleration due to gravity m/s^2          9.732.  9.796.      1.669  1.789        3.749.  3.728


Average g value for location m/s^2               9.764                  1.729                   3.739

answered by: Tulsiram Garg

> hi, thanks for the answer. I appreciate it.

Genius360 Sat, Nov 6, 2021 2:32 PM

> please give an explanation. I need it urgently like in the next 5 minutes

Genius360 Mon, Nov 8, 2021 1:21 AM